OpenStudy (anonymous):

Algebra 1 - Solving Rational Expressions. 3 + 1 = 6 --- ------ -- m-4 3(m-4) m Please show your work so I can understand how you got your answer.

OpenStudy (chmvijay):

WHAT IS lcm CAN U TELL

OpenStudy (anonymous):

What?

OpenStudy (chmvijay):

9+1/3(M-4) =6/M 10 =18(M-4)/M 10m=18m-72 72=18M-10M 72=8M M=72/8 m =9

OpenStudy (anonymous):

Thank you

OpenStudy (chmvijay):

UR WELCOME :)